OpenStudy (anonymous):

can someone please help me answer this question? Solve the system using elimination. 2x +6y = -12 5x - 5y =10 a. (2,1) b. (0,-2) c. (-2,0) d. (1,2)

OpenStudy (anonymous):

Choose one of the variable to eliminate let's go with 'x' We need to make the x terms in each equation so that when we add them they cancel. The LCD of 2 and 5 is 10. 2x+6y=-12 multiply by (-5) 5x-5y=10 multiply by (2) (-10x-30y = 60) (eqn 3) +(10x-10y=20 ) (eqn 4) Add equation 3 to equation 4 above you get -40y = 80 solve for y: y=-2 Go back to the original equations: choose equation (1)or (2) and substitute y=-2 2x+6(-2)=-12 solve for x = 0 Therefore, your answer as an ordered pair is (0,-2) which corresponds to b.

OpenStudy (anonymous):

Do you have any questions on what I did?